### v1

@@ 4.4 Network Architectures and Training
We use a UNet [Ronneberger et al., 2015, Song et al., 2021] to train both the score and consistency networks, with four down- and upsampling layers. For the four layers, we use convolutions with 128, 128, 256, and 256 channels, respectively, and $3\times 3$ kernels, SiLU activations, group normalization, and an attention layer at the architecture bottleneck.
We train the score network with the ADAM optimizer [Kingma and Ba, 2015] for 200 epochs, with a batch size of $1$, a learning rate of $2e^{-4}$, and an exponential moving average (EMA) over the model weights with a decay rate of 0.999 (see SI for more details).
The CM model is trained for 200 epochs following [Song et al., 2023], with the RADAM optimizer [Liu et al., 2021] and the same batch size, learning rate, and EMA schedule as the score network.
@@ 4.5 Scale-Consistent Downscaling
As shown in [Rissanen et al., 2023, Bischoff and Deck, 2023], adding Gaussian noise with a chosen variance to an image (or fluid dynamical snapshot) results in removing spatial patterns up to a specific spatial scale associated with the amount of added noise. The trained generative model can then replace the noise with spatial patterns learned from the training data up to the chosen spatial scale.
In principle, the spatial scale can be chosen depending on the given downscaling task, e.g., ESM resolution or variable. In general, ESM fields are too smooth at high small spatial scales, which presents a key problem for Earth system modelling in general and impact assessments in particular. More specifically, when comparing the frequency distribution of spatial precipitation fields in terms of spatial power spectral densities (PSDs), it can be seen that ESMs lack high-frequency spatial variability, or spatial intermittency, that is a key characteristic of precipitation [Hess et al., 2022]. Hence, a natural choice for the spatial scale to be preserved in the ESM fields is the intersection of the PSDs from the ESM and the ground truth ERA5 [Bischoff and Deck, 2023] (see Fig. 3), i.e., the scale where the ESM fields become too smooth.
For Gaussian noise, the variance as a function of time $t$ can be related to the PSD of a given wavenumber $k$ and the grid size $N$ by [Bischoff and Deck, 2023]
@@
The diffusion bridge (DB) [Bischoff and Deck, 2023] starts with the forward SDE in Eq. 1, initialized with a precipitation field from the POEM ESM. The forward SDE is then integrated until $t=t^{*}$. The reverse SDE (Eq. 2), initialized at $t=t^{*}$, then denoises the field again, adding structure from the target ERA5 distribution.
For the CM approach, at inference we apply the “stroke guidance” technique [Meng et al., 2022, Song et al., 2023], where we first sample a noised ESM field $\mathbf{\tilde{x}}^{\mathrm{ESM}}\in\mathbb{R}^{d}$ with variance corresponding to $t^{*}$,
@@
thus highly efficiently generating realistic samples $\mathbf{\hat{x}}$ that preserve spatial patterns of the ESM up to scale $k^{*}$.

### v3

@@ 4.4 Network Architectures and Training
We use a 2D UNet [ronneberger_u-net_2015, song_score-based_2021] from the Diffusers library to train both the score and consistency networks from scratch, with four down- and upsampling layers. For the four layers, we use convolutions with 128, 128, 256, and 256 channels, respectively, and $3\times 3$ kernels, SiLU activations, group normalization, and an attention layer at the architecture bottleneck. The network has, in total, around 27M trainable parameters.
We train the score network with the ADAM optimizer [kingma_adam_2015] for 200 epochs, with a batch size of $1$, a learning rate of $2e^{-4}$, and an exponential moving average (EMA) over the model weights with a decay rate of 0.999 (see SI for more details).
The CM model is trained for 150 epochs following [song_consistency_2023], with the RADAM optimizer [liu_variance_2021] and the same batch size, learning rate, and EMA schedule (with an initial decay rate of $\mu_{0}=0.9$) as the score network. We find that the loss decreases in a stable way throughout the training (Fig. S1). The training of 150 epochs takes around six and a half days for the CM and four and a half days for the SDE on a NVIDIA V100 32GB GPU. A summary of the training hyperparameters is given in the SI in Tab. S1.
@@ 4.5 Scale-Consistent Downscaling
As shown in [rissanen_generative_2023, bischoff_unpaired_2024], adding Gaussian noise with a chosen variance to an image (or fluid dynamical snapshot) results in removing spatial patterns up to a specific spatial scale associated with the amount of added noise. The trained generative model can then replace the noise with spatial patterns learned from the training data up to the chosen spatial scale.
In principle, the spatial scale can be chosen depending on the given downscaling task, e.g. related to the ESM resolution or variable. Hence, our method allows for much more flexibility after training, where the optimal spatial scale could be defined with respect to any given metric. In general, ESM fields are too smooth at small spatial scales, which presents a key problem for Earth system modelling in general and impact assessments in particular. More specifically, when comparing the frequency distribution of spatial precipitation fields in terms of spatial power spectral densities (PSDs), it can be seen that ESMs lack the high-frequency spatial variability, or spatial intermittency, that is a key characteristic of precipitation [hess_physically_2022]. Hence, a natural choice for the spatial scale to be preserved in the ESM fields is the intersection of the PSDs from the ESM and the ground truth ERA5 [bischoff_unpaired_2024] (see Fig. 3), i.e., the scale where the ESM fields become too smooth.
For Gaussian noise, the variance as a function of time $t$ can be related to the PSD of a given wavenumber $k$ and the grid size $N$ by [bischoff_unpaired_2024]
@@
The diffusion bridge (DB) [bischoff_unpaired_2024] starts with the forward SDE in Eq. 1, initialized with a precipitation field from the POEM ESM. The forward SDE is then integrated until $t=t^{*}$. The reverse SDE (Eq. 2), initialized at $t=t^{*}$, then denoises the field again, adding structure from the target ERA5 distribution.
For the CM approach, at inference we apply the “stroke guidance” technique [meng_sdedit_2022, song_consistency_2023], where we first sample a noised ESM field $\mathbf{\tilde{x}}^{\mathrm{ESM}}\in\mathbb{R}^{d}$ with variance corresponding to $t^{*}$,
@@
thus highly efficiently generating realistic samples $\mathbf{\hat{x}}$ that preserve unbiased spatial patterns of the ESM up to scale $k^{*}$.
Data Availability
The ERA5 reanalysis data is available for download at the Copernicus Climate Change Service (C3S) (https://cds.climate.copernicus.eu/datasets/reanalysis-era5-single-levels?tab=download). The simulation data from the POEM ESM is available at https://doi.org/10.5281/zenodo.4683086 [data-gmd-2020-436].
Code Availability
The Python code for processing and analyzing the data, together with the PyTorch [paszke_pytorch_2019] code for training [code-natmi-2024], will be made available on GitHub: https://github.com/p-hss/consistency-climate-downscaling.git, and CodeOcean https://doi.org/10.24433/CO.2150269.v1[codeocean-natmi-2024].
Acknowledgments
The authors would like to thank the three anonymous referees for their valuable remarks and suggestions. The authors thank Katherine Deck and Milan Klöwer for their valuable input. NB and PH acknowledge funding by the Volkswagen Foundation, as well as the European Regional Development Fund (ERDF), the German Federal Ministry of Education and Research, and the Land Brandenburg for supporting this project by providing resources on the high-performance computer system at the Potsdam Institute for Climate Impact Research. MA acknowledges funding under the Excellence Strategy of the Federal Government and the Länder through the TUM Innovation Network EarthCare. This is ClimTip contribution #X; the ClimTip project has received funding from the European Union’s Horizon Europe research and innovation programme under grant agreement No. 101137601. BP acknowledges funding by the National Key R&D Program of China (Grant NoS. 2023YFC3007700).
Author contributions
PH and NB conceived the research and designed the study with input from MA and BP. PH performed the model training and numerical analysis. PH, MA, BP and NB interpreted and discussed the results. PH wrote the manuscript with input from MA, BP and NB.
Competing interests
The authors declare no competing interests.
[…]